# Optimization Insights into Deep Diagonal Linear Networks

### v2

@@ 2.1 On the behavior of the flow
From equation 5, we can prove the following properties on the trajectory of the layers.
@@
    $\forall t\geqslant 0,\leavevmode\nobreak\ \frac{d}{dt}\left(u^{j}\odot u^{j}\right)(t)=\frac{d}{dt}\left(u^{k}\odot u^{k}\right)(t)=-\theta(t)\odot\nabla L(\theta(t)).$ (6)
@@
    $\forall t\geqslant 0,\leavevmode\nobreak\ u^{j}(t)^{\odot 2}-u^{j}(0)^{\odot 2}=u^{k}(t)^{\odot 2}-u^{k}(0)^{\odot 2}.$ (7)
@@
Before commenting the previous result, we introduce the main assumption on the initialization of the network that we will consider in the next:
$\forall i\in[d],\ \leavevmode\nobreak\ \#\left(\operatorname*{arg\,min}_{k\in[L]}|u^{k}_{i}(0)|\right)=1.$ ($\mathcal{A}$)
The preceding condition indicates that, for each component $i\in[d]$, the node with the minimal absolute value is unique across all layers. For instance, initializing each layer with the same values violates this assumption; while setting them randomly is almost surely valid.
It should be noted that this initialization assumption is relatively weak in comparison to those presented in the existing literature, even when considering only two layers. For Diagonal Linear Networks, [Woodworth et al., 2020] impose a layer to be zero before training while in the context of deep matrix factorization, [Arora et al., 2019] impose the layers to be balanced i.e.
@@
One can notice that by assuming $\mathcal{A}$, the second claim of Proposition 1 ensures that for each component $i$, only one layer $j$ can cross zero (and this layer is the one with the minimal absolute value at initialization). This phenomenon can be observed in Figure 3 where for the first component, only the third layer changes its sign throughout the training process. This layer is that with minimal absolute value at initialization.
@@
We use this feature to rewrite the dynamic of the layers.
@@
Suppose that assumption $\mathcal{A}$ is satisfied. Then, it allows us to define $\left(v^{j}\right)_{j\in[L]}\in\left(\mathbb{R}^{d}\right)^{L}$ as a permutation of $\left(u^{j}\right)_{j\in[L]}$ in the following way: for any $i\in[d]$,
$v^{1}_{i}=u^{j}_{i}\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ j=\arg\min_{k\in[L]}|u^{j}_{i}(0)|,$
@@
$v^{j}_{i}=\left\{\begin{aligned} &u^{1}_{i}\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=\arg\min_{k\in[L]}|u^{j}_{i}(0)|.\\ &u^{j}_{i}\leavevmode\nobreak\ \text{else.}\end{aligned}\right.$
In other words, $\left(v^{j}\right)_{j\in[L]}$ is a permutation of $\left(u^{j}\right)_{j\in[L]}$ where the first layer gathers all minimal absolute values at initialization. Note that we have $\theta=\bigodot_{j=1}^{L}v^{j}$ and it is trivial to show that the equations and properties stated above still hold for $\left(v^{j}\right)_{j\in[L]}$.
@@
$\forall t\geqslant 0,\leavevmode\nobreak\ \text{sign}\left(v^{j}(t)\right)=\text{sign}\left(v^{j}(0)\right),$
where $\text{sign}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ returns the sign for every component. By using the first claim of Proposition 1, we get that
$\forall j\in\llbracket 2,L\rrbracket,\leavevmode\nobreak\ \forall t\geqslant 0,\leavevmode\nobreak\ v^{j}(t)=\text{sign}\left(v^{j}(0)\right)\odot\sqrt{v^{1}(t)^{\odot 2}+v^{j}(0)^{\odot 2}-v^{1}(0)^{\odot 2}}.$
@@
$\forall t\geqslant 0,\leavevmode\nobreak\ \frac{dv^{1}(t)}{dt}+\text{sign}\left(\bigodot_{j\neq 1}v^{j}(0)\right)\odot\left(\bigodot_{j\neq 1}\sqrt{v^{1}(t)^{\odot 2}+\Delta_{j}}\right)\odot\nabla L(\theta(t))=0,$ (8)
@@
From that point, it appears natural to apply the strategy proposed by [Woodworth et al., 2020] to study Diagonal Linear Networks: the authors leverage a formulation of $\theta$ to derive an implicit bias which is also linked to a Mirror Flow (for further details, see Appendix A.1; for another example to which it applies, see Appendix A.3). However, if $L>2$, equation 8 does not allow us to get an analytical expression of $v^{1}(t)$ and thus of $\theta(t)$. As a consequence, this approach cannot be applied in this situation. Note that in the case $L=2$, equation 8 gives for any $t\geqslant 0$,
@@
with $\Delta=v^{2}(0)^{\odot 2}-v^{1}(0)^{\odot 2}$, allowing to compute $\theta$ and recover the analysis of [Woodworth et al., 2020].
2.2 It is actually a Mirror Flow
Since this setting does not fit the framework of [Woodworth et al., 2020], we base our analysis on the work of [Li et al., 2022] to prove the following theorem. Although our analysis demonstrates that $\theta$ is the solution of a Mirror Flow, it does not yield the associated entropy function (nor the implicit bias).
@@ 2.2 It is actually a Mirror Flow
Let $\theta\in\mathbb{R}^{d}$ be parameterized as $\theta=\bigodot_{j=1}^{L}u^{j}$ and $\left(u^{j}\right)_{j\in[L]}$ follow the dynamic described in equation 5. If the initialization of the network satisfies $\mathcal{A}$, then $\theta$ follows a Mirror Flow dynamic, i.e there exists a convex entropy function $\mathcal{Q}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that:
@@
Proof.
The proof of this claim consists in showing that the reparameterization induced by Deep Diagonal Linear Networks satisfies the hypotheses necessary to the application of [Li et al., 2022, Theorem 4.8]. It is first necessary to introduce some notation.
@@
$\forall n\in[Ld],\leavevmode\nobreak\ {\bm{u}}_{i}=u^{(i-1)\%L+1}_{\lfloor(i-1)/L\rfloor+1},$
@@
[…]
@@
$\forall w\in\mathbb{R}^{L\times d},\leavevmode\nobreak\ \forall i\in[d],\leavevmode\nobreak\ G_{i}(w)=\prod_{j=(i-1)L+1}^{iL}w_{j}.$ (10)
Showing the desired claim then requires to prove that $G:\mathcal{M}\rightarrow\mathbb{R}^{d}$ is a commuting and regular parameterization for a well-chosen smooth manifold $\mathcal{M}$ of $\mathbb{R}^{L\times d}$. Then, under an additional technical assumption that we discuss and if ${\bm{u}}(0)\in\mathcal{M}$, [Li et al., 2022, Theorem 4.8] ensures that the trajectory of $\theta$ is the solution of a Mirror Flow. The detailed computations are given in Appendix A.4.∎
In the previous theorem we have shown that, under mild assumptions on the inizialization, running Gradient Flow on the reparametrization $(u^{j})_{j\in[L]}$ given by a Deep Diagonal Linear Network generates a trajectory in $\theta(t)=\odot_{j=1}^{L}u^{j}(t)$ that is the solution of a Mirror Flow on the function $\mathcal{L}$ with entropy $\mathcal{Q}$. The strategy adopted does not lead to an explicit expression for the entropy $\mathcal{Q}$ (and so for the implicit bias induced by the reparametrization). On the other hand, we can guarantee that the entropy $\mathcal{Q}$ is convex and so, for instance, that the quantity $\mathcal{L}(\theta(t))$ is non-increasing along time.
Note that [Yun et al., 2021] provide a unifying analysis of the implicit bias of neural networks allowing to derive the entropy $\mathcal{Q}$ for Deep Diagonal Linear Networks for a particular initialization of the layers (see Corollary 5). It remains an open question whether the tensor formulation introduced in this paper can be extended to accommodate the general assumption A, which could be an interesting direction for future research.
@@ 2.3 Convergence guarantees
We now give additional convergence results on the trajectory of $\theta$ under the assumption $\mathcal{A}$. First, notice that $\theta$ is a solution of the following dynamical system
@@
Since we assume that assumption $\mathcal{A}$ holds, the second claim of Proposition 1 ensures that for any $i\in[d]$, if $j\notin\arg\min_{k\in[L]}|u^{k}_{i}(0)|$, then $|u^{j}_{i}(t)|>0$ for all $t\geqslant 0$. More precisely, using the first claim of Proposition 1, we obtain that $u^{j}_{i}(t)^{2}\geqslant u^{j}_{i}(0)^{2}-u^{l}_{i}(0)^{2}$ where $l=\arg\min_{k\in[L]}|u^{k}_{i}(0)|$. Therefore, for any $i\in[d]$,
$\forall t\geqslant 0,\leavevmode\nobreak\ \sum_{j=1}^{L}\left(\prod_{k\neq j}u^{k}_{i}(t)^{2}\right)$ $=\sum_{j=1,\leavevmode\nobreak\ j\neq l}^{L}\left(\prod_{k\neq j}u^{k}_{i}(t)^{2}\right)+\prod_{k\neq l}u^{k}_{i}(t)^{2}$ (13)
@@
As a consequence, we get the following properties:
@@
Thus, we get the following result which applies, for example, if $\mathcal{L}$ is the squared loss function.
@@
where we use the lower bound $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$. is define in equation 14. Supposing that $\mathcal{L}$ satisfies the Polyak-Łojasiewicz inequality, there exists $\mu>0$ such that:
@@
This bound on the suboptimality gap reveals that the initialization of the network significantly affects the convergence rate of its training. In this regard, it is erroneous to assume that the linear rate necessarily implies rapid convergence, as the parameter $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$ plays an important role. If the gap between the minimal absolute value of a component across layers and the others is small, then $\sigma$ is also considerably small, which affects the speed of convergence during training. Figure 4 emphasizes this point, showing that larger values at initialization lead to a better convergence rate.
Figure 4: Evolution of $\log\left(\mathcal{L}(\theta(t))-\mathcal{L}^{*}\right)$ according to time for three $6$-layer networks with different initialization. The loss function is defined as $\mathcal{L}:\theta\mapsto\|X\theta-y\|^{2}$ with $X\in\mathbb{R}^{10\times 8}$ and $y\in\mathbb{R}^{10}$ generated randomly. Each network is initialized with a first layer having components equal to 00. The initial value of the remaining layers of the first network (in blue) is generated randomly, while that of the second (in orange) and the third (in green) are respectively equal to $1.4$ and $1.8$ times the values of the firs component-wise.
This phenomenon has been documented in the literature for similar models (e.g. see [Chou et al., 2023a, Chou et al., 2024]) and may indicate that initializing the network with larger values would be beneficial. However, it is also well known that the initialization of a network should be chosen small enough to induce sparsity of the approximated solutions and consequently generalize well, see [Gunasekar et al., 2017, Arora et al., 2019, Woodworth et al., 2020, Moroshko et al., 2020, Chizat et al., 2019]. To address the challenge of slow training, [Chou et al., 2024] suggest employing weight normalization, a reparameterization technique that promotes comparable sparsity in the solution while ensuring reasonable convergence rates.

### v3

@@ 2.1 On the behavior of the flow
From equation 3, we can prove the following properties on the trajectory of the layers weights.
@@
    $\forall t\geqslant 0,\penalty 10000\ \frac{d}{dt}\left(u^{j}\odot u^{j}\right)(t)=\frac{d}{dt}\left(u^{k}\odot u^{k}\right)(t)=-\theta(t)\odot\nabla L(\theta(t)).$ (4)
@@
    $\forall t\geqslant 0,\penalty 10000\ u^{j}(t)^{\odot 2}-u^{j}(0)^{\odot 2}=u^{k}(t)^{\odot 2}-u^{k}(0)^{\odot 2}.$ (5)
@@
The following assumption on the initialization of the network is considered in the following and useful to interpret the above result. Assume that,
$\forall i\in[d],\ \penalty 10000\ \#\left(\operatorname*{arg\,min}_{k\in[L]}|u^{k}_{i}(0)|\right)=1.$ ($\mathcal{A}$)
The above condition states that, for each component $i\in[d]$, the node with the minimal absolute value is unique across all layers. For instance, initializing each layer with the same values violates this assumption; while setting them randomly is almost surely valid. This choice of the initialization is weak compared to other in the literature, even when considering only two layers. For shallow diagonal linear networks, [37] impose a layer to be zero before training while in the context of deep matrix factorization, [6] impose the layers to be balanced i.e.
@@
Back to Proposition 1, we note that, by Assumption $\mathcal{A}$, the second claim ensures that for each component $i$, only one layer $j$ can cross zero (and this layer is the one with the minimal absolute value at initialization). This phenomenon can be observed in Figure 3 where for the first component, only the third layer changes its sign throughout the training process. This layer is the one with minimal absolute value at initialization.
@@
We use this observation to rewrite the dynamic of the layers.
@@
Assumption $\mathcal{A}$ allows to define $\left(v^{j}\right)_{j\in[L]}\in\left(\mathbb{R}^{d}\right)^{L}$ as a permutation of $\left(u^{j}\right)_{j\in[L]}$ in the following way: for any $i\in[d]$,
$v^{1}_{i}=u^{j}_{i}\penalty 10000\ \text{where}\penalty 10000\ j=\arg\min_{k\in[L]}|u^{k}_{i}(0)|,$
@@
$v^{j}_{i}=\left\{\begin{aligned} &u^{1}_{i}\penalty 10000\ \text{if}\penalty 10000\ j=\arg\min_{k\in[L]}|u^{k}_{i}(0)|.\\ &u^{j}_{i}\penalty 10000\ \text{else.}\end{aligned}\right$.
In other words, $\left(v^{j}\right)_{j\in[L]}$ is a permutation of $\left(u^{j}\right)_{j\in[L]}$ where the first layer gathers all minimal absolute values at initialization. Note that we have $\theta=\bigodot_{j=1}^{L}v^{j}$ and it is trivial to show that the properties discussed above for $\left(u^{j}\right)_{j\in[L]}$ still hold for $\left(v^{j}\right)_{j\in[L]}$.
@@
$\forall t\geqslant 0,\penalty 10000\ \text{sign}\left(v^{j}(t)\right)=\text{sign}\left(v^{j}(0)\right),$
where $\text{sign}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ returns the sign for every component. By the first claim of Proposition 1, we get that
$\forall j\in\llbracket 2,L\rrbracket,\penalty 10000\ \forall t\geqslant 0,\penalty 10000\ v^{j}(t)=\text{sign}\left(v^{j}(0)\right)\odot\sqrt{v^{1}(t)^{\odot 2}+v^{j}(0)^{\odot 2}-v^{1}(0)^{\odot 2}}.$
@@
$\forall t\geqslant 0,\penalty 10000\ \frac{dv^{1}(t)}{dt}+\text{sign}\left(\bigodot_{j\neq 1}v^{j}(0)\right)\odot\left(\bigodot_{j\neq 1}\sqrt{v^{1}(t)^{\odot 2}+\Delta_{j}}\right)\odot\nabla L(\theta(t))=0,$ (6)
@@
The above observations could suggest to consider the strategy proposed by [37] to study shallow Diagonal Linear Networks. There, a suitable expression for $\theta$ is leveraged to derive an implicit bias linked to a Mirror Flow (for further details, see Appendix A.1; for another example to which it applies, see Appendix A.3). However, if $L>2$, equation 6 does not allow us to get an analytical expression of $v^{1}(t)$ and thus of $\theta(t)$. As a consequence, this approach cannot be applied in this situation. Note that in the case $L=2$, equation 6 gives for any $t\geqslant 0$,
@@
with $\Delta=v^{2}(0)^{\odot 2}-v^{1}(0)^{\odot 2}$, allowing to compute $\theta$ and recover the analysis of [37].
2.2 Mirror Flow for Deep Diagonal Linear Network
Since our setting does not fit the framework of [37], we base our analysis on the approach proposed in[21] to prove the following theorem.
@@ 2.2 Mirror Flow for Deep Diagonal Linear Network
Let $\theta\in\mathbb{R}^{d}$ be parameterized as in equartion equation 1 with $\left(u^{j}\right)_{j\in[L]}$ following the dynamic in equation 3. If the initialization of the network satisfies $\mathcal{A}$, then $\theta$ follows a Mirror Flow dynamic, i.e. there exists a convex Legendre function $\mathcal{Q}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that:
@@
In particular,
$\mathcal{L}(\theta(t))-\mathcal{L}^{*}\leq\frac{D_{Q}(\theta^{*},\theta_{0})}{t},$ (8)
where $\theta^{*}\in\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}(\theta)$ and $D_{Q}(\theta^{*},\theta)=Q(\theta^{*})-Q(\theta)-\langle\nabla Q(\theta),\theta^{*}-\theta\rangle$.
Proof. The proof of this claim consists in showing that the reparameterization induced by Deep Diagonal Linear Networks satisfies the hypotheses necessary to the application of [21, Theorem 4.8]. It is first necessary to introduce some notation.
@@
$\forall n\in[Ld],\penalty 10000\ {\bm{u}}_{i}=u^{(i-1)\%L+1}_{\lfloor(i-1)/L\rfloor+1},$
@@
${\bm{u}}=\left(u^{1}_{1}\penalty 10000\ u^{2}_{1}\penalty 10000\ \dots\penalty 10000\ u^{L}_{1}\penalty 10000\ u^{1}_{2}\penalty 10000\ \dots\dots\dots\penalty 10000\ u^{1}_{d}\penalty 10000\ \dots\penalty 10000\ u^{L}_{d}\right)^{T}.$
@@
$\forall w\in\mathbb{R}^{L\times d},\penalty 10000\ \forall i\in[d],\penalty 10000\ G_{i}(w)=\prod_{j=(i-1)L+1}^{iL}w_{j}.$ (9)
Showing the desired claim then requires to prove that $G:\mathcal{M}\rightarrow\mathbb{R}^{d}$ is a commuting and regular parameterization for a well-chosen smooth manifold $\mathcal{M}$ of $\mathbb{R}^{L\times d}$. Then, under an additional technical assumption that we discuss and if ${\bm{u}}(0)\in\mathcal{M}$, [21, Theorem 4.8] ensures that the trajectory of $\theta$ is the solution of a Mirror Flow. The detailed computations are given in Appendix A.4.
The above results shows that, under mild assumptions on the inizialization, Gradient Flow on the reparametrization $(u^{j})_{j\in[L]}$ generates a trajectory in the effective parameters $\theta(t)=\odot_{j=1}^{L}u^{j}(t)$ that is the solution of a Mirror Flow on the function $\mathcal{L}$ with entropy $\mathcal{Q}$. Our approach ensure that $\mathcal{Q}$ is convex and so, for instance, that the quantity $\mathcal{L}(\theta(t))$ is non-increasing along time. Indeed, in the following quantitative convergence guarantees will also be derived. However, our proof strategy does not lead to an explicit expression for the entropy $\mathcal{Q}$ (and hence for the implicit bias induced by the reparametrization). Interestingly, the results in [39] allow to derive the expression of the entropy $\mathcal{Q}$ for a particular initialization, namely when $u^{L}(0)=0$ and $u^{j}(0)=\alpha\bar{u}$ for every $j\in[L-1]$, with $\alpha>0$ and $\bar{u}\in\mathbb{R}^{d}$ satisfying $\bar{u}_{i}\neq 0$ for all $i\in[d]$. Indeed, under this initialization, and when $\mathcal{L}$ is the least squares loss, the trajectory $\theta(t)$ converges to $\theta_{\infty}$, a global minimizer. Moreover, the limiting solution minimizes a distance that behaves like the $\ell_{1}$-norm as $\alpha^{L}\to 0$, and like the $\ell_{2}$-norm as $\alpha^{L}\to\infty$. For further details, we refer the interested reader to [39, Corollary 5]. It remains unclear whether the formulation introduced in this paper can be extended to accommodate the choice of initialization as in assumption $\mathcal{A}$. This question is left as a direction for future research. Instead, we next derive suitable convergence guarantees.
@@ 2.3 Convergence guarantees
We derive additional convergence results on the trajectory of $\theta$ under assumption $\mathcal{A}$. First, note that $\theta$ is a solution of the following dynamical system
@@
Since we assumption $\mathcal{A}$ holds, the second claim of Proposition 1 ensures that for any $i\in[d]$, if $j\notin\arg\min_{k\in[L]}|u^{k}_{i}(0)|$, then $|u^{j}_{i}(t)|>0$ for all $t\geqslant 0$. More precisely, using the first claim of Proposition 1, we obtain that $u^{j}_{i}(t)^{2}\geqslant u^{j}_{i}(0)^{2}-u^{l}_{i}(0)^{2}$ where $l=\arg\min_{k\in[L]}|u^{k}_{i}(0)|$. Therefore, for any $i\in[d]$,
$\forall t\geqslant 0,\penalty 10000\ \sum_{j=1}^{L}\left(\prod_{k\neq j}u^{k}_{i}(t)^{2}\right)$ $=\sum_{j=1,\penalty 10000\ j\neq l}^{L}\left(\prod_{k\neq j}u^{k}_{i}(t)^{2}\right)+\prod_{k\neq l}u^{k}_{i}(t)^{2}$ (12)
@@
Then, we get the following properties:
@@
Then, we get the following result which applies, for instance, if $\mathcal{L}$ is the squared loss function.
@@
where we use the lower bound $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$ obtained in equation 13. Supposing that $\mathcal{L}$ satisfies the Polyak-Łojasiewicz inequality, there exists $\mu>0$ such that:
@@
The above bound on the suboptimality gap reveals that the initialization of the network significantly affects the convergence rate of its training and that $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$ plays an important role. Consider for example the case where $\mathcal{L}$ is $\mu$ strongly convex and hence satisfies the $\mu$ PL inequality. In this setting, vanilla gradient flow applied to the effective parameters gives a rate $e^{-2\mu t}(\mathcal{L}(\theta_{0})-\mathcal{L}^{*})$. Instead, considering the overparameterization equation 1 we obtain in the new variables a problem which is non convex in general. Despite this, the previous result shows that gradient flow on such non convex problem may result in the original variable $\theta$ in a faster converegence rate, as long as the initialization is such that $\sigma>1$.
This becomes particularly clear when considering the initialization considered in [39].
Corollary 1.
Let $\theta\in\mathbb{R}^{d}$ be parameterized as in equation equation 1 and $\left(u^{j}\right)_{j\in[L]}$ follow the dynamic in equation 3. Suppose that $\arg\min_{\theta\in\mathbb{R}^{d}}\mathcal{L}(\theta)\neq\emptyset$ and let $\mathcal{L}^{*}$ denote $\min_{\theta\in\mathbb{R}^{d}}\mathcal{L}(\theta)$. If the initialization is such that $u^{L}(0)=0$ and $u^{j}(0)=\alpha\bar{u}$ for every $j\in[L-1]$, with $\alpha>0$ and $\bar{u}\in\mathbb{R}^{d}$ satisfying $|\bar{u}_{i}|^{2}\geqslant\lambda>0$ for all $i\in[d]$, and the loss function $\mathcal{L}$ satisfies a Polyak-Łojasiewicz property with parameter $\mu>0$, then for all $t\geqslant 0$,
$\mathcal{L}(\theta(t))-\mathcal{L}^{*}\leqslant e^{-2\sigma\mu t}\left(\mathcal{L}(\theta(0))-\mathcal{L}^{*}\right),\qquad\sigma=(\alpha^{2}\lambda)^{L-1}.$
Proof.
The initialization of the network directly ensures that assumption $\mathcal{A}$ is satisfied and the explicit computation of $\sigma$ is straightforward.∎
In this case, if $\alpha^{2}\lambda>1$ then $\sigma$ is increasing as a function of the number of layers, and this emphasizes the beneficial effect of deep overparametrizations on the convergence rate. Figure 4 hihglights the impact of the initial amplitude of the weights, showing that larger values at initialization lead to a better convergence rate.
Figure 4: Evolution of $\log\left(\mathcal{L}(\theta(t))-\mathcal{L}^{*}\right)$ according to time for three 66-layer networks with different initialization. The loss function is defined as $\mathcal{L}:\theta\mapsto\|X\theta-y\|^{2}$ with $X\in\mathbb{R}^{10\times 8}$ and $y\in\mathbb{R}^{10}$ generated randomly. Each network is initialized with a first layer having components equal to 0. The initial value of the remaining layers of the first network (in blue) is generated randomly, while that of the second (in orange) and the third (in green) are respectively equal to $1.4$ and $1.8$ times the values of the first component-wise.
The above results highlights a phenomenon documented in the literature for similar models (e.g. see [10, 13]) and may indicate that initializing the network with large values would be beneficial. However, it is also well known that the initialization of a network could be chosen small enough to induce sparsity of the approximated solutions, see [18, 6, 37, 26, 9]. Indeed, to address the challenge of slow training, [13] suggest employing weight normalization, a reparameterization technique that promotes comparable sparsity in the solution while ensuring reasonable convergence rates.